# Quantum Algorithm for Solving the Advection Equation using Hamiltonian Simulation

### v2

@@ I Introduction
Quantum algorithms for solving PDEs can be delineated into two categories: fully quantum algorithms that implement quantum circuits to evolve the quantum state as described by the PDE of interest, and quantum-classical hybrid algorithms where a quantum computer is used for a specific task in a larger, classical computation. Fully quantum approaches generally excel at solving linear PDEs because quantum operators act linearly on quantum superpositions, allowing algorithms based on the finite difference method (FDM) [10, 11, 12, 13], the finite element method (FEM) [14, 15] and spectral methods [16] to be effectively represented quantum mechanically. Encoding the solution from $N=2^{n}$ grid points within the amplitudes $\alpha_{j}$ of an $n$-qubit quantum state |$\ket{\psi}=\sum_{j=1}^{N}\alpha_{j}\ket{j}$ leads to an exponentially growing capacity to store information and an inherent quantum parallelism when processing it. Amplitude encoding does not allow for the inspection of the full solution as with classical methods, but rather the extraction of global statistics into the limited output space. This may be adequate depending on the context and make previously intractable problems tractable. Fully quantum approaches are not limited to linear PDEs as techniques have been proposed based on the derivation of the nonlinear Schrödinger equation using mean-field techniques [17] and Carleman linearisation [18] to tackle nonlinear PDEs, but these are generally limited to weakly nonlinear interactions. On the other hand, variational quantum algorithms (VQAs) [19] for solving optimisation problems have been used as the basis for hybrid algorithms that have demonstrated a greater capability of tackling nonlinear PDEs [20, 21, 22]. Kyriienko et al. [21] used a machine learning strategy where differentiable quantum circuits were trained to solve nonlinear differential equations. Jaksch et al. [22] extended a quantum algorithm for solving nonlinear problems [20] to fluid dynamics, evaluating cost functions from matrix product state representations of the flow [23] to obtain a polynomial upper bound on the depth of the variational network. VQA-based algorithms are of interest because their potential to operate on near-term hardware, [19] though a definitive quantum advantage is yet to be demonstrated.
@@
Quantum algorithms for solving PDEs that do not depend on a QLSA have also been developed. Costa et al. [11] put forward a quantum algorithm that evolves a quantum state according to the wave equation using Hamiltonian simulation without the need for a QLSA, aside from in the proposed generation of the initial conditions. A practical implementation of this algorithm was then developed and analysed by Suau et al. [33], confirming that the gate requirements agreed with the theoretical complexity. In a different study, Budinski [34] proposed a quantum algorithm for the advection-diffusion equation centred around the lattice Boltzmann method (LBM) that tracks the evolution of particle distribution functions on a grid. The main challenge of implementing the non-unitary collision step of the LBM was achieved with the LCU method. Simulating ODEs in the form of Eq. (2) as Hamiltonian simulations has received much recent attention [35, 36, 37, 38, 32]. An et al. [38] expressed the evolution as a linear combination of Hamiltonian simulation (LCHS) problems, which is a special case of the LCU method. The algorithm was later improved upon [32] by further generalising the method and leading to the discovery of faster-decaying coefficients in the weighted sum, improving the $\epsilon$ dependency. Jin et al. [35, 36, 37] provided an alternative viewpoint on this problem by introducing a Schrödingerisation method that maps linear PDEs to a higher-dimensional system of Schrödinger equations, then solving with Hamiltonian simulation. Both the Schrödingerisation [35, 36, 37] and the LCHS [38, 32] algorithms assume $M$ to be negative semi-definite, applicable to problems and numerical schemes that maintain or dampen the solution, but do not amplify it. This applies to upwind and central finite difference schemes for discretising the advection equation, but not downwind schemes. A time-marching algorithm for the ODE problem was proposed by Fang et al. [39] by explicitly integrating the PDE over short time steps, as is a common practice in the classical numerical solution of differential equations. This operator is not unitary so has a probability of failure in a block-encoding strategy, ordinarily leading to an exponentially decaying success probability in the simulation time $T$. This was overcome by applying a uniform singular value amplification at each time step, resulting in a runtime with a quadratic dependence on $T$. The absence of a QLSA in these methods generally leads to improved state preparation costs [32] and favourable poly-logarithmic scaling in the allowable error per time step using sparse Hamiltonian simulation [40] or LCU [30] algorithms. Other methods of quantum matrix multiplication with the potential to be applied to solving PDEs using explicit time advancement were compared by Shao [41] based on the swap test, singular value estimation and HHL algorithms. However, all of these methods utilise QPE as a subroutine, thus limiting their $\epsilon$ dependence and the number of state initialisation queries required.
@@ II The Algorithm
for spatial grid point $m$ and temporal location $t$. Equation (3) can be solved for $\phi_{m}^{t+1}$, obtaining an equation to advance the solution in time. A vector $\vec{\phi}_{t}=[\phi_{0}^{t},\phi_{1}^{t},\dots,\phi_{N-1}^{t}]$ can be constructed from $\phi$ to write Eq. (3) as a matrix transformation
@@
While divergence-free advection is a norm-preserving process, the truncation errors from the discretisation procedure in Eq. (3) lead to the matrix $A$ being non-unitary, i.e. $AA^{\dagger}\neq I$. Therefore, to enact $A$ on a quantum state |$\ket{\phi_{t}}$, the non-unitary Hamiltonian embedding procedure described by Gingrich and Williams [43] is used. A Hamiltonian $H$ is constructed from $A$ in the form
@@
$=\exp\left(\begin{bmatrix}0&A\theta\\ -A^{\dagger}\theta&0\end{bmatrix}\right)$ (7)
where the symbol $\theta$ is the Hamiltonian evolution time for a single time step. The time step size is encoded in the Hamiltonian, so $\theta$ affects the accuracy of a time step and its probability of success rather than the evolution time itself. The exponential function produces a block matrix with the structure
$\Omega=\begin{bmatrix}\widetilde{I}&\widetilde{A}\\ \widetilde{B}&\widetilde{I}\end{bmatrix}$ (8)
and when applied to a solution register |$\ket{\phi_{t}}$ supplemented by an ancilla qubit initialised as |$\ket{1}$, it produces the state
@@
For the advection equation, the matrices $\widetilde{A}$ and $\widetilde{I}$ retain the relative structure of $A$ and the identity matrix $I$ respectively, and hence, a subspace of the Hilbert space evolves as described by the advection equation. Postselecting the ancilla qubit in the state |$\ket{0}$ collapses the solution register to |$\ket{\phi_{t+1}}=\widetilde{A}\ket{\phi_{t}}$, and if postselection fails the state collapses to $\widetilde{I}\ket{\phi_{t}}\approx\ket{\phi_{t}}$ allowing the state to be reused for further computation. The quantum circuit representations of a successful and unsuccessful time step are shown in Fig. 1.
@@ III.3 Overall Error Bound
Figure 3: Surface plot of the error bound per nondimensional simulation time as a function of $r$ and $\theta$ for (a) the advection equation and (b) the heat equation. For the heat equation in (b), the surface has been truncated at $\epsilon/T=20$ since $\epsilon$ asymptotically approaches infinity for $r_{h}\to 0.25$ and $r_{h}\to 0.5$.
The overall error bound for the algorithm can be evaluated using $\epsilon=N_{T}||E_{A}||+N_{F}||E_{I}||$, where $N_{T}$ is the required number of successful time steps and $N_{F}$ is the expected number of failed time steps. Since the number of time steps required is inversely proportional to the time step, $N_{T}=T/r$, where $T$ is the simulation time in nondimensional form and $N_{F}=N_{T}/P_{\text{min}}-N_{T}$. Combining these expressions, the error per the simulation time evaluates to
@@ IV Complexity Analysis
The number of qubits grows as $n=O(\log N)$ since the computational grid is compressed into the amplitudes of the exponentially growing computational basis states. The circuit depth grows linearly with the required number of time steps $N_{T}\propto TN^{1/D}$. The number of time steps growing linearly with the desired simulation time $T$ is intuitive, but the $N^{1/D}$ factor with the number of grid points $N$ requires a further breakdown. It is assumed that a $D$-dimensional domain of equal side lengths is discretised by $N=N_{x}^{D}$ grid points, where $N_{x}=N_{y}=N_{z}$ for $D>1$. The CFL condition $r_{\text{max}}=\text{max}(u_{j})\Delta t/\Delta x<1$ constrains the time step size to the grid spacing, where halving $\Delta x$ requires halving $\Delta t$, or equivalently, doubling $N_{x}$ requires doubling $N_{T}$. Given that $N_{T}\propto N_{x}$ and $N=N_{x}^{D}$, then $N\propto N_{T}^{D}$, or inversely $N_{T}\propto N^{1/D}$.
@@
Figure 4: Evolution of the quantum amplitudes from statevector simulations of a 2D laminar channel flow described by the advection equation at different time intervals for (a) $\theta=\pi/(1+\sqrt{r_{\text{max}}^{2}+1})$ while varying $r_{\text{max}}$ and (b) $r_{\text{max}}=0.25$ while varying $\theta$.
@@ V.1 Laminar Channel Flow
The velocity field $\vec{u}=[u,v]$ is described by the analytical solution to the Navier-Stokes equations in this configuration, known as a plane Poiseuille flow, which has a parabolic profile of $u$ leading to the nondimensional CFL parameter to be defined as
@@ VI Conclusions
A quantum algorithm for solving the advection equation, a linear PDE prevalent in various scientific and engineering industries, was presented. The algorithm uses sparse Hamiltonian simulation to embed the discrete time-marching operator $A$ into the Hamiltonian $H$, resulting in a unitary operator $\Omega=e^{-iH\theta}$ that encodes $A$ to a high accuracy regardless of the values of $\theta$. This ensures that a time step is completed with high accuracy and probability. Postselection failure does not require further state initialisation queries since the resulting operation closely approximates the identity matrix, having a minimal impact on the quantum state and allowing the computation to continue. The algorithm applies to multidimensional problems with arbitrary boundary conditions and finite difference stencils.
@@
The typical advantages of the proposed approach over other algorithms are that the runtime grows linearly with the simulation time, the state can be reused on postselection failure requiring a single copy of the initial state, any combination of finite difference stencils can be used, and for its simplicity of implementation. Other algorithms that can be applied to the advection equation are typically aimed at solving homogeneous ODEs in the form of Eq. (2), which the advection equation reduces to when discretised in space. In comparison to QLSA-based algorithms [24, 25, 28, 27], the proposed approach excels due to the single copy of the initial quantum state required. The optimal QLSA [31] requires $O(\kappa\log(1/\epsilon))$ queries to the state preparation oracle, which can become prohibitive when pursuing a practical quantum advantage. The quantum time-marching algorithm [39] is conceptually the most similar to the algorithm presented here, although results in a runtime with a quadratic dependence on $T$ compared to a linear dependence in the present algorithm. Schrödingerisation [35, 36, 37] and LCHS [38, 32] algorithms both assume that the real part of the coefficient matrix $M$ is negative semi-definite, which applies to central and upwind finite difference schemes for the advection equation. Occasionally, however, the use of downwind schemes cannot be avoided, such as when resolving the flow near a computational boundary. Therefore, the flexibility of the proposed approach and its simplicity of implementation make it a preferred choice.

### v3

@@ I Introduction
Quantum algorithms for solving PDEs can be delineated into two categories: fully quantum algorithms that implement quantum circuits to evolve the quantum state as described by the PDE of interest, and quantum-classical hybrid algorithms where a quantum computer is used for a specific task in a larger, classical computation. Fully quantum approaches generally excel at solving linear PDEs because quantum operators act linearly on quantum superpositions, allowing algorithms based on the finite difference method (FDM) [10, 11, 12, 13], the finite element method (FEM) [14, 15] and spectral methods [16] to be effectively represented quantum mechanically. Encoding the solution from $N=2^{n}$ grid points within the amplitudes $\alpha_{j}$ of an $n$-qubit quantum state |$\ket{\psi}=\sum_{j=1}^{N}\alpha_{j}\ket{j}$ leads to an exponentially growing capacity to store information and an inherent quantum parallelism when processing it. Amplitude encoding does not allow for the inspection of the full solution as with classical methods, but rather the extraction of global statistics into the limited output space. This may be adequate depending on the context and make previously intractable problems tractable. Fully quantum approaches are not limited to linear PDEs as techniques have been proposed based on the derivation of the nonlinear Schrödinger equation using mean-field techniques [17] and Carleman linearisation [18] to tackle nonlinear PDEs, but these are generally limited to weakly nonlinear interactions. On the other hand, variational quantum algorithms (VQAs) [19] for solving optimisation problems have been used as the basis for hybrid algorithms that have demonstrated a greater capability of tackling nonlinear PDEs [20, 21, 22]. Kyriienko et al. [21] used a machine learning strategy where differentiable quantum circuits were trained to solve nonlinear differential equations. Jaksch et al. [22] extended a quantum algorithm for solving nonlinear problems [20] to fluid dynamics, evaluating cost functions from matrix product state representations of the flow [23] to obtain a polynomial upper bound on the depth of the variational network. VQA-based algorithms are of interest because of their potential to operate on near-term hardware, [19] though a definitive quantum advantage is yet to be demonstrated.
@@
Quantum algorithms for solving PDEs that do not depend on a QLSA have also been developed. Costa et al. [11] put forward a quantum algorithm that evolves a quantum state according to the wave equation using Hamiltonian simulation without the need for a QLSA, aside from in the proposed generation of the initial conditions. A practical implementation of this algorithm was then developed and analysed by Suau et al. [33], confirming that the gate requirements agreed with the theoretical complexity. In a different study, Budinski [34] proposed a quantum algorithm for the advection-diffusion equation centred around the lattice Boltzmann method (LBM) that tracks the evolution of particle distribution functions on a grid. The main challenge of implementing the non-unitary collision step of the LBM was achieved with the LCU method. Simulating ODEs in the form of Eq. (2) as Hamiltonian simulations has received much recent attention [35, 36, 37, 38, 32]. An et al. [38] expressed the evolution as a linear combination of Hamiltonian simulation (LCHS) problems, which is a special case of the LCU method. The algorithm was later improved upon [32] by further generalising the method and leading to the discovery of faster-decaying coefficients in the weighted sum, improving the $\epsilon$ dependency. Jin et al. [35, 36, 37] provided an alternative viewpoint on this problem by introducing a Schrödingerisation method that maps linear PDEs to a higher-dimensional system of Schrödinger equations, then solving with Hamiltonian simulation. Both the Schrödingerisation [35, 36, 37] and the LCHS [38, 32] algorithms assume the Hermitian part of $M$ to be negative semi-definite, applicable to problems and numerical schemes that maintain or dampen the solution, but do not amplify it. This applies to upwind and central finite difference schemes for discretising the advection equation, but not downwind schemes. A time-marching algorithm for the ODE problem was proposed by Fang et al. [39] by explicitly integrating the PDE over short time steps, as is a common practice in the classical numerical solution of differential equations. This operator is not unitary so has a probability of failure in a block-encoding strategy, ordinarily leading to an exponentially decaying success probability in the simulation time $T$. This was overcome by applying a uniform singular value amplification at each time step, resulting in a runtime with a quadratic dependence on $T$. The absence of a QLSA in these methods generally leads to improved state preparation costs [32] and favourable poly-logarithmic scaling in the allowable error per time step using sparse Hamiltonian simulation [40] or LCU [30] algorithms. Other methods of quantum matrix multiplication with the potential to be applied to solving PDEs using explicit time advancement were compared by Shao [41] based on the swap test, singular value estimation and HHL algorithms. However, all of these methods utilise QPE as a subroutine, thus limiting their $\epsilon$ dependence and the number of state initialisation queries required.
@@ II The Algorithm
for spatial grid point $m$ and temporal location $t$. When solved classically, the forward-time central-space scheme in Eq. (3) is unstable for hyperbolic PDEs such as the advection equation, requiring upwind schemes for conditional stability [42]. The quantum representation of such forward-time schemes in the present algorithm is stable for arbitrary finite difference stencils, as will be demonstrated later in this section. Equation (3) can be solved for $\phi_{m}^{t+1}$, obtaining an equation to advance the solution in time. A vector $\vec{\phi}_{t}=[\phi_{0}^{t},\phi_{1}^{t},\dots,\phi_{N-1}^{t}]$ can be constructed from $\phi$ to write Eq. (3) as a matrix transformation
@@
While divergence-free advection is a norm-preserving process, the truncation errors from the discretisation procedure in Eq. (3) lead to the matrix $A$ being non-unitary, i.e. $A^{\dagger}A\neq I$ and $AA^{\dagger}\neq I$. Therefore, to enact $A$ on a quantum state |$\ket{\phi_{t}}$, the non-unitary Hamiltonian embedding procedure described by Gingrich and Williams [44] is used. A Hamiltonian $H$ is constructed from $A$ in the form
@@
$=\exp\begin{bmatrix}0&A\theta\\ -A^{\dagger}\theta&0\end{bmatrix}$ (7)
where the symbol $\theta$ is the Hamiltonian evolution time for a single time step. The time step size is encoded in the Hamiltonian, so $\theta$ affects the accuracy of a time step and its probability of success rather than the evolution time itself. The exponential function produces a block matrix with the structure [44]
$\exp\begin{bmatrix}0&X\\ Y&0\end{bmatrix}$ $=\begin{bmatrix}\cosh(\sqrt{XY})&X\frac{\sinh(\sqrt{YX})}{\sqrt{YX}}\\ Y\frac{\sinh(\sqrt{XY})}{\sqrt{XY}}&\cosh(\sqrt{YX})\end{bmatrix}$
$=\begin{bmatrix}\cos(\theta\sqrt{AA^{\dagger}})&A\frac{\sin(\theta\sqrt{A^{\dagger}A})}{\sqrt{A^{\dagger}A}}\\ -A^{\dagger}\frac{\sin(\theta\sqrt{AA^{\dagger}})}{\sqrt{AA^{\dagger}}}&\cos(\theta\sqrt{A^{\dagger}A})\end{bmatrix}$ (8)
where the top-right and top-left blocks of Eq. (8) are termed $\widetilde{A}$ and $\widetilde{I}$ respectively. The $\widetilde{A}$ matrix is closely proportional to $A$ for small values of $\theta$ or when $\sqrt{A^{\dagger}A}\approx I$ since $\sin(I\theta)=I\sin(\theta)$, leading to $\widetilde{A}\approx A\sin(\theta)$. Furthermore, the $\widetilde{I}$ matrix is closely proportional to $I$ for small values of $\theta$ or when $\sqrt{AA^{\dagger}}\approx I$, leading to $\widetilde{I}\approx I\cos(\theta)$. For the advection equation evolution operator in Eq. (5), $\sqrt{A^{\dagger}A}$ and $\sqrt{AA^{\dagger}}$ can be approximated by $I+O(r^{2})$ for small values of $r$. Equation (8) also reveals that the singular values $\sigma_{i}(\widetilde{A})=\sin(\sigma_{i}(A)\theta)$ are bounded by 1, stabilising the scheme regardless of the finite difference stencil. Applying $\Omega$ to a solution register |$\ket{\phi_{t}}$ supplemented by an ancilla qubit initialised as |$\ket{1}$ produces the state
@@
Postselecting the ancilla qubit in the state |$\ket{0}$ collapses the solution register to |$\ket{\phi_{t+1}}=\widetilde{A}\ket{\phi_{t}}$, and if postselection fails, the state collapses to $\widetilde{I}\ket{\phi_{t}}\approx\ket{\phi_{t}}$, allowing the state to be reused for further computation. The quantum circuit representations of a successful and unsuccessful time step are shown in Fig. 1.
@@ III.3 Overall Error Bound
Figure 3: Surface plot of the error bound per non-dimensional simulation time as a function of $r$ and $\theta$ for (a) the advection equation and (b) the heat equation. For the heat equation in (b), the surface has been truncated at $\epsilon/T=20$ since $\epsilon$ asymptotically approaches infinity for $r_{h}\to 0.25$ and $r_{h}\to 0.5$.
The overall error bound for the algorithm can be evaluated using $\epsilon=N_{T}||E_{A}||+N_{F}||E_{I}||$, where $N_{T}$ is the required number of successful time steps and $N_{F}$ is the expected number of failed time steps. Since the number of time steps required is inversely proportional to the time step, $N_{T}=T/r$, where $T$ is the simulation time in non-dimensional form and $N_{F}=N_{T}/P_{\text{min}}-N_{T}$. Combining these expressions, the error per the simulation time evaluates to
@@ IV Complexity Analysis
The number of qubits grows as $n=O(\log N)$ since the computational grid is compressed into the amplitudes of the exponentially growing computational basis states. The circuit depth grows linearly with the required number of time steps $N_{T}\propto TN^{1/D}$. The number of time steps growing linearly with the desired simulation time $T$ is intuitive, but the $N^{1/D}$ factor with the number of grid points $N$ requires a further breakdown. It is assumed that a $D$-dimensional domain of equal side lengths is discretised by $N=N_{x}^{D}$ grid points, where $N_{x}=N_{y}=N_{z}$ for $D>1$. The CFL number $r_{\text{max}}=\text{max}(u_{j})\Delta t/\Delta x$ constrains the time step size to the grid spacing, where halving $\Delta x$ requires halving $\Delta t$, or equivalently, doubling $N_{x}$ requires doubling $N_{T}$. Given that $N_{T}\propto N_{x}$ and $N=N_{x}^{D}$, then $N\propto N_{T}^{D}$, or inversely $N_{T}\propto N^{1/D}$.
@@
Figure 4: Evolution of the quantum amplitudes from statevector simulations of a 2D laminar channel flow described by the advection equation at different time intervals for (a) $\theta=\pi/(1+\sqrt{r_{\text{max}}^{2}+1})$ while varying $r_{\text{max}}$ and (b) $r_{\text{max}}=0.25$ while varying $\theta$. The simulations use a 2nd-order-accurate central finite difference stencil.
@@ V.1 Laminar Channel Flow
The velocity field $\vec{u}=[u,v]$ is described by the analytical solution to the Navier-Stokes equations in this configuration, known as a plane Poiseuille flow, which has a parabolic profile of $u$ leading to the non-dimensional CFL parameter to be defined as
@@ VI Conclusions
A quantum algorithm for solving the advection equation, a linear PDE prevalent in various scientific and engineering industries, was presented. The algorithm uses sparse Hamiltonian simulation to embed the discrete time-marching operator $A$ into the Hamiltonian $H$, resulting in a unitary operator $\Omega=e^{-iH\theta}$ that encodes $A$ to a high accuracy regardless of the values of $\theta$. Postselection failure does not require further state initialisation queries since the resulting operation closely approximates the identity matrix, having a minimal impact on the quantum state and allowing the computation to continue. The algorithm applies to multidimensional problems with arbitrary boundary conditions and finite difference stencils.
@@
The typical advantages of the proposed approach over other algorithms are that the runtime grows linearly with the simulation time, the state can be reused on postselection failure requiring a single copy of the initial state, any combination of finite difference stencils can be used, and for its simplicity of implementation. Other algorithms that can be applied to the advection equation are typically aimed at solving homogeneous ODEs in the form of Eq. (2), which the advection equation reduces to when discretised in space. In comparison to QLSA-based algorithms [24, 25, 28, 27], the proposed approach excels due to the single copy of the initial quantum state required. The optimal QLSA [31] requires $O(\kappa\log(1/\epsilon))$ queries to the state preparation oracle, which can become prohibitive when pursuing a practical quantum advantage. The quantum time-marching algorithm [39] is conceptually the most similar to the algorithm presented here, although results in a runtime with a quadratic dependence on $T$ compared to a linear dependence in the present algorithm. Schrödingerisation [35, 36, 37] and LCHS [38, 32] algorithms both assume that the Hermitian part of the coefficient matrix $M$ is negative semi-definite, which applies to central and upwind finite difference schemes for the advection equation. Occasionally, however, the use of downwind schemes cannot be avoided, such as when resolving the flow near a computational boundary. Therefore, the flexibility of the proposed approach and its simplicity of implementation make it a preferred choice.